OpenStudy (anonymous):

sqrt9b^3 - sqrt25b^3+ sqrt49b^3

OpenStudy (hoblos):

sqrt(9b^3) = sqrt(3^2 * b^2 * b ) = (3b)sqrt(b)

OpenStudy (hoblos):

can you try doing the same for sqrt25b^3 and sqrt49b^3 ??

OpenStudy (anonymous):

ok

OpenStudy (hoblos):

tell me what you get so that i can check it for you ;)

OpenStudy (anonymous):

5sqrt5b

OpenStudy (hoblos):

sqrt25b^3 = (5b)sqrt(b)

OpenStudy (hoblos):

do you understand how im doing that?

OpenStudy (anonymous):

ok

OpenStudy (hoblos):

so what is sqrt49b^3 ?